A train P start from Mokama at 5:00 pm reached Hazipur at 6:00 pm. An another train Q starts from Hazipur at 5:00 pm and reaches Mokama at 6:30 pm. At what time, two trains will cross each other ?

Grade:10

1 Answers

Arun
4 years ago
Let the distance between Mokama and Hazipur be D km  
Time taken by train P to cover D km = 1 h 
Time taken by Q to cover D km = 3/2 h
∴ Speed of train P = D km/h 
Speed of train Q = 2D/3 km/h .
Let they cross each other T h after 5 : 00 pm. 
Then, DT + 2DT/ 3 = D 
⇒ T + 2T/3 = 1  
⇒ T(1 + 2/3) = 1 
⇒ T = 3/5h = (3/5) x 60 = 3 x 12 = 36 min.

Hence, the two trains meet 36 min after 5 : 00 pm. 
∴ Meeting time = 5:36 pm